OpenStudy (some.random.cool.kid):

What is the simplified form of left parenthesis 27 x to the power of negative six end power y to the power of 12 end power right parenthesis to the power of start fraction two over three end fraction end power

OpenStudy (some.random.cool.kid):

@zepdrix

OpenStudy (some.random.cool.kid):

@triciaal

OpenStudy (some.random.cool.kid):

@nincompoop @whpalmer4 @pooja195 @paki @ganeshie8

OpenStudy (some.random.cool.kid):

mind helping?

zepdrix (zepdrix):

\[\large\rm \left[abc\right]^{n}=a^n b^n c^n\]When we have a bunch of stuff being raised to a power, we have to apply the exponent to each of them.

zepdrix (zepdrix):

\[\large\rm \left[27x^{-6}y^{12}\right]^{2/3}=27^{2/3}(x^{-6})^{2/3}(y^{12})^{2/3}\]Ok with that step? :)

OpenStudy (some.random.cool.kid):

mm got it am i supposed to add across? for the powers part?

OpenStudy (some.random.cool.kid):

or exponents i should say?

zepdrix (zepdrix):

Have to apply one of your exponent rules from that point:\[\large\rm \color{orangered}{(x^a)^b=x^{a\cdot b}}\]When you have a power and another power, you `multiply`

OpenStudy (some.random.cool.kid):

mmm ok ...

zepdrix (zepdrix):

So for the exponent on x, we'll multiply -6 and 2/3.

OpenStudy (some.random.cool.kid):

-6 x 12? I belive. what each of those two thirds?

OpenStudy (some.random.cool.kid):

oh

OpenStudy (some.random.cool.kid):

I belive that is -4? correct...

zepdrix (zepdrix):

\[\large\rm \left[27x^{-6}y^{12}\right]^{2/3}=27^{2/3}x^{-4}(y^{12})^{2/3}\]k looks good :)

OpenStudy (some.random.cool.kid):

mmm next step is ? -4 x 12 or -4 x the other 2/3rds?

OpenStudy (some.random.cool.kid):

or is it -4 x 12 and 2/3 x 2/3?

OpenStudy (some.random.cool.kid):

or is it just multiply across lol? i believe its what i said first though

zepdrix (zepdrix):

Notice that these two things have `different bases`. One has a base of x, the other a base of y, that's letting us know that they won't interact with one another. \(\large\rm x^{a}\cdot x^{b}=x^{a+b}\) This rule that you're familiar with only works `when they have the same base`.

zepdrix (zepdrix):

So the -4 on the x isn't going to interact in any way with the powers on y.

OpenStudy (some.random.cool.kid):

right so find the lcm?

zepdrix (zepdrix):

You have \(\large\rm (y^{12})^{2/3}\). Our exponent rule tells us to `multiply` these powers, ya? 12 * 2/3?

OpenStudy (some.random.cool.kid):

8

zepdrix (zepdrix):

\[\large\rm \left[27x^{-6}y^{12}\right]^{2/3}=\color{orangered}{27^{2/3}}x^{-4}y^{8}\]k good :) almost done

OpenStudy (some.random.cool.kid):

is this going to be 8 over 4?

OpenStudy (some.random.cool.kid):

no

zepdrix (zepdrix):

Try to pay attention to what I've been saying. `They have different bases`, so they won't interact with one another in any way.

OpenStudy (some.random.cool.kid):

its not nvm i answered my own thing i belive you leave those alon now right so now i have 9 x^4y^8?

zepdrix (zepdrix):

yay good job \c:/

OpenStudy (some.random.cool.kid):

:D mind helping me with a few more?

zepdrix (zepdrix):

stay cool kid (⌐▨_▨) oh sure

OpenStudy (some.random.cool.kid):

Ok great I have another one this one is a bit different then the last a little bit ...

OpenStudy (some.random.cool.kid):

A student uses the quadratic formula to solve a quadratic equation and determines that one of the solutions is x equals negative five plus square root of negative 147 end square root. What are the values of a and b if this solution is written in the form x = a + bi, where a and b are real numbers?

OpenStudy (some.random.cool.kid):

hold on i need to attach the numbers part this is too messy

zepdrix (zepdrix):

\[\large\rm x=-5+\sqrt{-147}\]

OpenStudy (some.random.cool.kid):

there we go :D

zepdrix (zepdrix):

Well, recall that we define the imaginary unit i, to be \(\large\rm \sqrt{-1}\)

zepdrix (zepdrix):

So we can do something clever with the square root in our problem.

zepdrix (zepdrix):

\[\large\rm \sqrt{-147}=\quad\sqrt{-1\cdot147}=\quad\sqrt{-1}\cdot\sqrt{147}\]I broke it up in a clever way to pull the sqrt(-1) outside like that. Do you see the i?

OpenStudy (some.random.cool.kid):

yep I have used this method before...

zepdrix (zepdrix):

\[\large\rm x=-5+\sqrt{-147}\]\[\large\rm x=-5+\color{orangered}{\sqrt{-1}}\cdot\sqrt{147}\]

OpenStudy (some.random.cool.kid):

first before you continue

OpenStudy (some.random.cool.kid):

shouldnt I be trying to find the sqrt?

OpenStudy (some.random.cool.kid):

and trying to figure out whetther its irrational or not?

OpenStudy (some.random.cool.kid):

because I believe that is 12.124355653

zepdrix (zepdrix):

Yes, try to simplify the square root if you're able to. In doesn't matter which order you do these steps in though :) No no, we don't want an ugly decimal approximation.

OpenStudy (some.random.cool.kid):

which is irrational so now you may preceed

OpenStudy (some.random.cool.kid):

eh 12.12

zepdrix (zepdrix):

no...

OpenStudy (some.random.cool.kid):

...

zepdrix (zepdrix):

You don't use your calculator for this step.

OpenStudy (some.random.cool.kid):

no i rounded lol

zepdrix (zepdrix):

But you first used a calculator.. that's not what we're doing here.

zepdrix (zepdrix):

You try to simplify the root by pulling out a `perfect square`. Example: \(\large\rm \sqrt{72}=\sqrt{2\cdot36}=\sqrt{2}\cdot\sqrt{36}=\sqrt{2}\cdot 6\)

OpenStudy (some.random.cool.kid):

oooo nvm its not 12.12

zepdrix (zepdrix):

So in my example here, I noticed that 72 was a multiple of 36, 36 can be taken out of the root since it's a perfect square. it comes out as a 6

OpenStudy (some.random.cool.kid):

its either 21 or 7sqrtof 3 I believe correct? not sure which one yet tough

OpenStudy (some.random.cool.kid):

though*

OpenStudy (some.random.cool.kid):

im not sure 6 is simplest of forms though?

zepdrix (zepdrix):

6sqrt2 is the simplified form of sqrt(72). That was just an example though.

zepdrix (zepdrix):

Break down 147, what are some factors that we can use? One little math trick is to add up the digits, 1 + 4 + 7 = 12. 12 is a multiple of 3, so this number is divisible by 3. So 147 = 3 * _

OpenStudy (some.random.cool.kid):

/uh

OpenStudy (some.random.cool.kid):

not sure

zepdrix (zepdrix):

Just use a calculator then :U 147 / 3

OpenStudy (some.random.cool.kid):

49

zepdrix (zepdrix):

ok\[\large\rm x=-5+\color{orangered}{\sqrt{-1}}\cdot\sqrt{147}\]\[\large\rm x=-5+\color{orangered}{\sqrt{-1}}\cdot\sqrt{3\cdot49}\]\[\large\rm x=-5+\color{orangered}{\sqrt{-1}}\cdot\sqrt{3}\cdot\sqrt{49}\]

zepdrix (zepdrix):

\[\large\rm x=-5+\color{orangered}{\mathcal i}\cdot\sqrt{3}\cdot\sqrt{49}\]What does sqrt(49) simplify to?

OpenStudy (some.random.cool.kid):

ok

OpenStudy (some.random.cool.kid):

7

zepdrix (zepdrix):

Good good good. Let's write the i in the back, so it looks a lil bit cleaner.\[\large\rm x=-5+7\sqrt3~\mathcal i\]

OpenStudy (some.random.cool.kid):

ok